OpenStudy (anonymous):

1. Simplify the following expression showing all steps and explain how the order of operations is applied in each step as the expression is simplified. Be sure to include the rules for addition, subtraction, multiplication, division, and the use of grouping symbols in the explanations of the steps used to simplify the expression. [(7 + 3) • 5 – 4] ÷ 2 + 2

OpenStudy (michele_laino):

here we have to do the addition first, so: what is 7+3=...?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

7+3=10

OpenStudy (anonymous):

oh oh lol sorry i taught you said 7-3, yeah its 10

OpenStudy (michele_laino):

ok! now we have to do the multiplication: 10 * 5=...?

OpenStudy (anonymous):

50

OpenStudy (michele_laino):

then we have to do the subtraction: 50-4=...?

OpenStudy (anonymous):

46

OpenStudy (michele_laino):

Now i divide by 2: 46:2=...?

OpenStudy (anonymous):

23

OpenStudy (michele_laino):

finally we have to do the addition: 23+2=...?

OpenStudy (anonymous):

25

OpenStudy (michele_laino):

correct! you got the result

OpenStudy (anonymous):

25?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

25 is the answer??

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

can you put everything step by step together like everything?

OpenStudy (michele_laino):

ok!: step #0 [(7 + 3) • 5 – 4] ÷ 2 + 2 step #1 [10 • 5 – 4] ÷ 2 + 2 step #2 [50 – 4] ÷ 2 + 2 step #3 46 ÷ 2 + 2 step #4 23 + 2 step #5 25

OpenStudy (anonymous):

thank you!!!!!!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

ok thanks for your help bye.